# Comment on “Source of black bounces in Rastall gravity”

### v1

@@ I Field Equations
where $L_{F}=\partial L/\partial F$, with the Lagrangian $L(F)$ being a general function of the electromagnetic scalar $F=F^{\mu\nu}F_{\mu\nu}/4$, $\phi$ is a scalar field, $V(\phi)$ is the potential related to the scalar field, $R$ is the trace of the Ricci tensor $R_{\mu}$, $g$ is the determinate of the metric $g_{\mu\nu}$, $\lambda$ is the Rastall parameter, $T^{\phi}_{\mu\nu}$ is the stress-energy tensor of the scalar field, and $T^{EM}_{\mu\nu}$ is the stress-energy tensor of the electromagnetic field. The stress-energy tensors are given by
@@ II.2 Bardeen-type
$\rho^{R}=\frac{2\gamma Q^{2}\left(\left(Q^{2}+r^{2}\right)^{3/2}-m\left(2Q^{2}+r^{2}\right)\right)}{\kappa^{2}\left(Q^{2}+r^{2}\right)^{7/2}}T=-p_{r}^{R}=-p_{t}^{R},$ (33)

### v2

@@ I Field Equations
where $L_{F}=\partial L/\partial F$, with the Lagrangian $L(F)$ being a general function of the electromagnetic scalar $F=F^{\mu\nu}F_{\mu\nu}/4$, $\phi$ is a scalar field, $V(\phi)$ is the potential related to the scalar field, $R$ is the trace of the Ricci tensor $R_{\mu\nu}$, $g$ is the determinate of the metric $g_{\mu\nu}$, $\lambda$ is the Rastall parameter, $T^{\phi}_{\mu\nu}$ is the stress-energy tensor of the scalar field, and $T^{EM}_{\mu\nu}$ is the stress-energy tensor of the electromagnetic field. The stress-energy tensors are given by
@@ II.2 Bardeen-type
$\rho^{R}=\frac{2\gamma Q^{2}\left(\left(Q^{2}+r^{2}\right)^{3/2}-m\left(2Q^{2}+r^{2}\right)\right)}{\kappa^{2}\left(Q^{2}+r^{2}\right)^{7/2}}=-p_{r}^{R}=-p_{t}^{R},$ (33)